a) The pH of an aqueous solution of 5.27×10-2 M ammonium perchlorate, NH4ClO4 (aq), is ________ .

This solution is _______ . (acidic, basic, or neutral)

b) The pH of an aqueous solution of 0.150 M sodium fluoride, NaF (aq), is ________ .

This solution is _______ . (acidic, basic, or neutral)
Answer #1

a)

concentration of NH4ClO4 = 5.27 x 10^-2 M

pKb of NH3 = 4.74

pH = 7 - 1/2(pKb + log C)

     = 7 - 1/2 (4.74 + log 0.0527)

pH = 5.27

This solution is acidic .

b)

concentration of NaF = 0.150

pKa of HF = 3.14

pH = 7 + 1/2 (pKa + log C)

     = 7 + 1/2 (3.14 + log 0.150)

pH = 8.16

This solution is basic